OpenStudy (dilla):

Simplify using method I or method II. m-4/m+4 + m-4/ m-2 / 1-2/m-2

OpenStudy (dilla):



OpenStudy (triciaal):

sorry the original is still not clear

OpenStudy (dilla):

m-4 divided by m+4 plus m-4 divided by m-2. All of that divided by 1- 2 divided by m-2

OpenStudy (loser66):

\[\dfrac{\dfrac{m-4}{m+4}+\dfrac{m+4}{m-2}}{1-\dfrac{2}{m-2}}\]